### v1

# Continuous Self-Improvement of Large Language Models by Test-time Training with Verifier-Driven Sample Selection

Abstract
Learning to adapt pretrained language models to unlabeled, out-of-distribution data is a critical challenge, as models often falter on structurally novel reasoning tasks even while excelling within their training distribution. We introduce a new framework called VDS-TTT - Verifier-Driven Sample Selection for Test-Time Training to efficiently address this. We use a learned verifier to score a pool of generated responses and select only from high ranking pseudo-labeled examples for fine-tuned adaptation. Specifically, for each input query our LLM generates $N$ candidate answers; the verifier assigns a reliability score to each, and the response with the highest confidence and above a fixed threshold is paired with its query for test-time training. We fine-tune only low-rank LoRA adapter parameters, ensuring adaptation efficiency and fast convergence. Our proposed self-supervised framework is the first to synthesize verifier driven test-time training data for continuous self-improvement of the model. Experiments across three diverse benchmarks and three state-of-the-art LLMs demonstrate that VDS-TTT yields up to a 32.29% relative improvement over the base model and a 6.66% gain compared to verifier-based methods without test-time training, highlighting its effectiveness and efficiency for on-the-fly large language model adaptation.
1 Introduction
Large-scale language models (LLMs) are typically developed under a fixed training-testing paradigm, where model parameters are learned once and remain unchanged during test time. However, a long-standing idea in machine learning advocates for test-time adaptation, i.e., updating the model using information specific to each test instance. This highlights the distinction between inductive learning, which aims to generalize from training data to unseen examples, and transductive learning, which enables models to dynamically adapt at test time. By embracing test-time updates, LLMs can better handle domain shifts and nonstationary environments ( Xiao and Snoek (2024)).
Test-time training (TTT) is a paradigm that revisits this idea, updating model parameters at test time based on each incoming test instance or batch, thereby performing on-the-fly transductive learning. Transductive or local learning methods, such as locally weighted regression and local fine-tuning—have been studied since the early days of machine learning (Bottou and Vapnik (1992); Cleveland (1979); Cleveland and Devlin (1988)), demonstrating the potential of instance-specific adaptation. TTT differs from continual pretraining. TTT adapts a pre-trained model during test time to each test instance using unsupervised objectives, requiring minimal or no labeled data and achieving rapid task-specific tuning. In contrast, Continual Pretraining (CPT) (Wu et al. (2024)) continuously updates a model on large, streaming datasets, supervised or unsupervised, aiming to gradually enrich its generalization capabilities across evolving domains.
A fundamental challenge in TTT is adapting to unseen test distributions without access to ground-truth labels. In the absence of labels, models cannot optimize the primary task loss directly and must instead depend on self-generated supervisory signals such as pseudo-labels or auxiliary self-supervised tasks such as in Meta Test-Time Training (MT3) ( Bartler et al. (2022)) or entropy-minimization objectives such as in TENT (Wang et al. (2020)). These surrogate signals, however, are often noisy and poorly aligned with the original task, leading to error accumulation and training instability (Wang et al. (2024)). Furthermore, adapting on all generated auxiliary samples imposes significant computational overhead and exacerbates catastrophic forgetting, when low-confidence or incorrect updates are applied uniformly during test time (Niu et al. (2022)).
Figure 1: In our VDS-TTT framework, we proceed through three sequential stages. First, Candidate Generation for Self-Annotation, where each input question $Q_{i}$ is passed to the pretrained LLM to produce a set of $N$ candidate responses $\{r_{1},...,r_{N}\}$. Second, Confidence-Guided Annotation, in which a verifier assigns a reliability score to each $r_{j}$, and we select the response $r^{*}$ only if its score exceeds a predefined threshold $\tau$, thereby forming the test-time training example $(Q_{i},r^{*})$. Finally, Test-Time Training, where we fine-tune the model by optimizing its parameters on the resulting pseudo-labeled dataset.
In this work, we introduce Verifier-Driven Sample Selection Test-Time Training (VDS-TTT), a new framework that synthesizes high confidence supervisory signals from unlabeled inputs. For each test example, VDS-TTT first generates a diverse set of candidate solutions by repetitively sampling from a large language model with a positive temperature. A pre-trained domain-specific verifier (for instance, a suite of unit tests in mathematical reasoning tasks) then scores each candidate on correctness. Test examples with all candidates receiving below acceptable confidence score from the verifier are not considered for use in training. For other test examples, candidates above a confidence threshold are retained as pseudo-labels to be used in training. In short, the model undergoes preference-based adaptation using only the verifier-approved samples (Figure 1). By leveraging a verifier to filter noisy outputs, VDS-TTT aligns optimization with the primary task objective, prevents catastrophic forgetting through modification of only LoRA adapter parameters, and self improves by dynamically adjusting to varying domain shifts across continuous test streams. Empirical results demonstrate that VDS-TTT achieves substantially robust adaptation compared to other approaches addressing label-scarce, real-world scenarios.
Our main contributions are as follows:
  • We introduce VDS-TTT, a new framework, a first to leverage a learned verifier to select a high confidence pseudo-label for use in supervised-finetuning at test time, addressing the absence of ground-truth labels during test time, and then fine-tunes low-rank LoRA adapter parameters to self-improve the base LLM on the fly.
  • Through comprehensive experiments across multiple benchmarks, we demonstrate that VDS-TTT consistently improves test-time accuracy compared to baseline without test time training, significantly outperforming methods that simply select samples based on verifier-computed confidence scores. This validates the effectiveness of integrating pseudo-label-based fine-tuning into the test time pipeline.
  • We show that increasing test time compute by increasing the number of test-time training iterations yields performance improvements, eventually matching or exceeding an oracle verifier at the first iteration, highlighting both the efficiency and robustness of our approach under heterogeneous solution distributions.
  • Experiments across three diverse benchmarks and three state-of-the-art LLMs demonstrate that VDS-TTT yields up to a 32.29% relative improvement over the base model and a 6.66% gain compared to verifier-based methods without test-time training.
2 Related Work
At test time, ground-truth labels are inherently unavailable, which prohibits the direct application of standard supervised objectives. Consequently, a rich line of work has emerged in the image and text domains around the development of self-supervised or unsupervised learning signals at the test time.
In the image domain, numerous self-supervised tasks have demonstrated effectiveness for this purpose, including rotation prediction (Sun et al. (2020)), meta-learning adaptations (Bartler et al. (2022)), and contrastive learning frameworks (Burns et al. (2023)). By tailoring these auxiliary losses to the current test distribution, such methods encourage the feature extractor to remain robust under domain shifts. However, applying these techniques in the text domain can alter the input semantics.
In the language modeling literature, different works tried to obtain unknown ground-truth labels to improve the model that can be categorized into extra-memory approaches through retrieval, internal or external feedback derived directly from model predictions, and reinforcement learning (RL) approaches.
For extra-memory approaches, retrieval-based solutions such as Hardt and Sun (2023) show positive gains in terms of perplexity across more than 20 downstream tasks. This is done by retrieving semantically similar contexts and performing a single gradient update on each of a handful of neighbors. Large retrieval indices incur significant storage and computational overhead: the full index can occupy between 810GB and 2.1TB. Moreover, naively selecting the top-N highest-scoring neighbors could often retrieve redundant or non-informative examples. To overcome this memory limitation, Hübotter et al. (2024) proposed a data-selection algorithm called SIFT that unifies retrieval with active learning. SIFT quantifies the model’s posterior uncertainty about a prompt after fine-tuning and selects examples with the least uncertainty. However, they still require memory to store information.
Another prominent category leverages internal or external feedback derived directly from model predictions. For example, Tent (Wang et al. (2020)) treats the prediction entropy as a proxy for uncertainty and fine-tunes the model during test time by minimizing this entropy. This approach did not consider model collapse into trivial constant output, which needs to be regularized (Press et al. (2024)).
Beyond fully unsupervised signals, certain task-specific frameworks exploit small amounts of human feedback or external reward functions as high-quality supervision (Ji et al. (2025)). These reward feedbacks are used in self-training approaches vi RL. ReST (Gulcehre et al. (2023)) and ReST-MCTS (Zhang et al. (2024)) are the most well-known RL-based approaches.
ReST Gulcehre et al. (2023) tries to formulate a growing-batch offline RL paradigm for LLMs but suffers from two issues. First is that it requires annotated data as during each Grow step, ReST samples new outputs from the current policy which, together with the original annotated data, are scored and retained only if their reward exceeds an iteratively increasing threshold. Second, the method uses a reduced learning rate across multiple Improve iterations to prevent overfitting, which could be better circumvented using parameter-efficient techniques.
Building on ReST, ReST-MCTS (Zhang et al. (2024)) targets reasoning-intensive tasks by embedding a Process Reward Model (PRM) within a Monte Carlo Tree Search framework to collect and value step-by-step reasoning traces. However, this technique has two major limitations. First, MCTS rollouts guided by a PRM usually require high computational cost that may not be easily applicable during test time. Second, the policy model requires high-quality full reasoning traces with known ground truth labels, which limits its direct application to test-time.
Most recently, Zuo et al. (2025) proposed an RL-based method that incorporates reward feedback during test-time. To derive a pseudo-label, this selects the answer with majority vote over a large ensemble of model generations and uses this answer for RL training. Specifically, Proximal Policy Optimization (PPO) (Schulman et al. (2017)) and Group Relative Policy Optimization (GRPO) (Shao et al. (2024)) optimize the policy at test-time. PPO follows a traditional actor–critic paradigm, generating one output per prompt, estimating advantages with a learned value network, and updating both policy and value networks with a clipped surrogate objective. In contrast, GRPO dispenses with a separate critic by sampling a large group of responses per prompt, computing group-normalized rewards to estimate advantages, and updating only the policy with surrogate clipping and a KL penalty. As mentioned in Zuo et al. (2025), RL-based methods typically depend on the pretrained distribution and are sensitive to the input difficulty level. This can be seen further in our experimental results section.
To alleviate the aforementioned challenges in the literature, we propose VDS-TTT, a continuous self-improvement paradigm that applies reward-based feedback at test-time. Our method has the following advantages: (1) does not require any labeled data; (2) simple and memory-efficient wihout extra memory requirements and expensive rollouts; (3) simpler to implement, offers greater stability in terms of convergence due to its low-variance gradient estimates Mukobi et al. (2023), and is more computationally efficient compared to existing approaches; (4) we update only low-rank LoRA adapter parameters rather than the entire model to be better suited to test-time that requires on-the-fly specialization while keeping model’s original performance since it is a lightweight adaptation so it can not dramtically change the base model.
3 Methodology
In this section, we introduce VDS-TTT (Algorithm 1), our new LLM-based, fully test-time training method. VDS-TTT operates in three sequential stages designed to improve model robustness at test time. First, the pretrained language model generates a diverse set of candidate solutions for each input problem, leveraging its capacity to explore multiple hypotheses. Next, a verifier-based scoring function evaluates each candidate. If all candidates are low scoring (below a pre-defined acceptable threshold), the input problem and its candidate solution(s) are not considered for use in further training. Otherwise, the highest-scoring output is selected as the provisional solution and paired with the input problem to form the sample data for training. The model is then fine-tuned on this sample data during testing, enabling it to adapt dynamically to distributional shifts and improve its performance on challenging inputs.
Algorithm 1 Verifier-Driven sample Selection for Test-Time Training (VDS-TTT)
1:Pretrained LLM $f_{\theta_{0}}$, verifier score function $s(\cdot,\cdot)$, temperature $T$, number of samples $N$, score threshold $\tau$, LoRA adapter steps $M$, learning rate $\eta$
2:Adapted adapter parameters $\Delta$
3:Initialize adapter $\Delta\leftarrow\mathbf{0}$
4:for all test query $q_{i}$ do $\triangleright$ Stage 1: Candidate Generation for Self-Annotation
5:     $\mathcal{R}\leftarrow\emptyset$
6:     for $j=1$ to $N$ do
7:         Sample response $r_{ij}\sim f_{\theta_{0}}\bigl{(}\cdot\mid q_{i};T\bigr{)}$triangleright temperature sampling
8:         Extract final answer $a_{ij}$ from $r_{ij}$
9:         $\mathcal{R}\leftarrow\mathcal{R}\cup\{(r_{ij},a_{ij})\}$
10:     end $\triangleright$ Stage 2: Confidence-Guided High-Quality Annotation
11:     Compute scores $s_{ij}\leftarrow s\bigl{(}r_{ij},a_{ij}\bigr{)}$ for each $(r_{ij},a_{ij})\in\mathcal{R}$
12:     Let $j^{*}=\arg\max_{1\leq j\leq N}s_{ij}$
13:     if $s_{ij^{*}}<\tau$ then
14:         continue $\triangleright$ skip low-confidence query
15:     end if
16:     Set pseudo-label $(r_{i},a_{i})\leftarrow(r_{ij^{*}},a_{ij^{*}})$
17:     Record score $s_{i}\leftarrow s_{ij^{*}}$triangleright Stage 3: Test-Time Training (LoRA Adaptation)
18:     for $m=1$ to $M$ do
19:         Compute loss
$\mathcal{L}(\Delta)=-\sum_{t=1}^{|r_{i}|}\log\,f_{\theta_{0}+\Delta}\bigl{(}r_{i,t}\mid q_{i},r_{i,<t}\bigr{)}$
20:         $\Delta\leftarrow\Delta-\eta\,\nabla_{\Delta}\,\mathcal{L}(\Delta)$
21:     end for
22:end for
23:return $\Delta$
3.1 Candidate Generation for Self-Annotation
In our approach, we automatically construct a labeled dataset $D=\{q_{i},r_{ij},a_{ij}\}_{j=1}^{N}$ without relying on human annotations, where $q_{i}$, $r_{ij}$, $a_{ij}$, and $N$ respectively denote input question, LLM generated detailed step by step responses, LLM generated corresponding final answer, and number of generated responses per input question. We do this by invoking the pretrained language model to generate for each question a diverse set of candidate answers, along with detailed step by step responses by employing temperature sampling, in which the softmax temperature $T$ is set at test time to control the randomness of the model’s output, thus letting the model explore more unconventional and less likely outputs (Renze (2024)).
3.2 Confidence-guided High-quality Data Annotation
At this stage, for each generated candidate response, we must determine which ones are suitable for inclusion in model training. Two paradigms exist: Verifier-Free methods (VF), which directly distill successful search or “thinking” traces without any external check, and Verifier-Based (VB) methods, which interleave a verification step to guide the search process. As demonstrated by Setlur et al. (2025), VF methods exhibit poor scalability when the base pretrained LLM produces a heterogeneous distribution over correct solution traces and is thus suboptimal, whereas VB methods maintain robust performance under these conditions. Consequently, we adopt a VB selection strategy to ensure both scalability and accuracy in assembling our test time training data.
For the $N$ candidate responses we apply a Best-of-N selection strategy: our pretrained verifier assigns a real-valued score to each generated response, and we choose the response $r^{*}=argmax_{y\in\{r_{1},...,r_{N}\}}s(r)$ with the highest score as the provisional answer. To ensure both accuracy and informativeness, we further impose a score threshold $\tau$: any response whose verifier score $s(r)$ falls below $\tau$ is discarded, so that only those candidates deemed sufficiently confident and high-quality are used to train the model, thus aligning better with the query task,. This two-step process, first selecting the single highest-scoring response and then filtering by a minimum confidence threshold, guarantees that our generated training data comprises diverse yet reliably correct solution traces.
3.3 Test Time Training
For each test query $q_{i}$ that is selected for TTT, we construct a test-time training set $D_{TTT}=\{q_{i},r_{i},a_{i},s_{i}\}$ where $s_{i}$ is the corresponding score provided by the Best-of-N verifier procedure. By treating each unlabeled response as labeled according to its confidence score, this method efficiently leverages unlabeled data without human annotation. Then, we fine-tune the pretrained LLM by minimizing the empirical risk:
$min_{\Delta}\sum_{(q,r,a,s)\in D_{TTT}}L(LM_{\Delta}(q,r),\hat{y})$ (1)
where $\Delta$ denotes the LoRA Hu et al. (2022) adapter parameters while the original model weights remain frozen. Freezing the base LLM and updating only LoRA adapter parameters preserves model stability, and avoids catastrophic forgetting while ensuring rapid convergence during test-time training. By leveraging pseudo-labels in a one-sample self-supervised manner, we enable on-the-fly model adjustment with each high confidence test instance, improving robustness under distributional shifts and enabling continuous self improvement.
4 Experiments
4.1 Settings
Datasets. In our experimental evaluation, we assess VDS-TTT on three established mathematical reasoning benchmarks: GSM8K, Math-500, and AIME1983-2024. GSM8K Cobbe et al. (2021) comprises 1319 grade-school math word problems, segmented test instance, each requiring 2–8 sequential elementary arithmetic operations to arrive at a final numeric answer. Math-500 is a curated subset of 500 problems drawn from the MATH dataset (12.5K competition-level questions with full, step-by-step solutions), Hendrycks et al. (2021), covering seven domains—Prealgebra, Algebra, Number Theory, Counting and Probability, Geometry, Intermediate Algebra, and Precalculus—as selected in Lightman et al. (2023). Finally, AIME1983-2024 comprises challenging problems from the American Invitational Mathematics Examination (AIME), collected over four decades (1983–2024). These competition-level problems require advanced reasoning, often combining algebraic manipulation with creative problem-solving and geometric insight. The American Mathematics Competitions (AMC) dataset contains 83 samples comprising a series of progressively challenging mathematical tests designed for middle and high school students. We report exact-match accuracy on held-out test splits, extracting final answers via regular-expression parsing and verifying equivalence with ground truth.
Baseline Models. To validate the generality of VDS-TTT, we apply our framework to four state-of-the-art LLMs: Llama-3.2-1B-Instruct, DeepSeek-R1-Distill-Qwen-1.5B, Llama-3.2-3B-Instruct, and LLaMA-3.1-8B-Instruct. Moreover, to reduce computational overhead, we adopt a lightweight reward model, Skywork-o1-Open-PRM-Qwen-2.5-1.5B.
4.2 Implementation details
In our experiments, we systematically vary the number of candidate responses $N\in\{2,4,8,16\}$ to evaluate its impact on Best-of-N selection performance. We adopt a stringent verifier confidence threshold $\tau=0.99$ for GSM8K and Math-500—mirroring pseudo-labeling best practices that apply high thresholds to filter out noisy labels, but relax $\tau$ to 0.9 for the more challenging AIME1983-2024 to prevent discarding the vast majority of low-scoring samples if $\tau=0.99$. Furthermore, we empirically concluded that updating only low-rank LoRA adapter parameters substantially outperforms full fine-tuning of deeper or randomly selected layers. Consequently, we integrate LoRA modules into all key projection components—q_proj, k_proj, v_proj, o_proj—as well as the MLP sublayers (mlp_gate_proj, mlp_up_proj, mlp_down_proj) of our base LLM. We set the LoRA rank to 128 in most cases where sufficient data is available for test-time training. For low-resource scenarios, such as AIME 2024, we reduce the LoRA rank to 8 to prevent overfitting and maintain training stability.
All experiments were performed on a machine equipped with an NVIDIA Tesla V100 PCIe GPU with 32GB of memory. As a representative example of total VDS-TTT execution time, applying the LLaMA-3.2-1B-Instruct model to the Math-500 dataset with 2, 4, 8, and 16 candidate responses resulted in total runtimes of approximately 1h30m, 2h10m, 3h6m, and 4h30m, respectively.
4.3 Results and Discussion
We first assess how VDS-TTT enhances model accuracy across different sampling budgets. Table 1 reports exact-match accuracy for our three baselines—Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, and DeepSeek-R1-Distill-Qwen-1.5B—on GSM8K, Math-500, and AIME1983–2024 using $N=2,4,8,16$ candidate responses. For each $N$, we compare three settings: (1) Base: the frozen model without any Test Time Training; (2) VB (Verifier-Based): where we generate $N$ samples and select the single highest-scoring response via our verifier; and (3) VDS-TTT: which fine-tunes LoRA adapters on all verifier-selected, high-confidence pseudo-labels.
The most dramatic improvements occur on AIME1983–2024, where Qwen-1.5B’s Base accuracy is near zero (0.54% for N=2), VB selection raises it to 0.88%, and just a single iteration of our VDS-TTT boosts it to 4.22%. With N=16, VDS-TTT accuracy increases to 6.96%. These results highlight VDS-TTT’s ability to extract reliable pseudo-labels and adapt models on-the-fly, even when the pretrained base model lacks prior competence in the test domain. Moreover, the gain between $N=2$ (2–6% over VB) and $N=4$ (1–5%) exceeds that between $N=8$ (1–3%) and $N=16$ (1–2%), indicating diminishing returns at larger sample sizes and suggesting that $N=4$ often suffices for robust adaptation. Overall, VDS-TTT not only amplifies verifier-based selection but also delivers robust, parameter-efficient adaptation under severe distribution shifts, with the greatest relative gains at smaller candidate budgets.
Table 1: Accuracy improvements across Math-500, GSM-8K, and AIME1983-2024 benchmarks for Llama3-1B, Qwen-1.5B, and Llama3-3B models using Base, Verifier-Based (VB), and VDS-TTT methods with varying numbers of samples ($N=2,4,8,16$). The results demonstrate the effectiveness of VDS-TTT in enhancing model performance, particularly in low-resource scenarios.
Model Name N Math-500 GSM-8K AIME1983-2024
Base VB VDS-TTT Base VB VDS-TTT Base VB VDS-TTT
Llama3-1B 2 20.8 24.00 28.00 40.18 52.31 55.88 3.42 6.22 8.37
4 26.80 30.60 59.67 62.40 6.43 9.29
8 33.60 36.60 63.08 63.84 9.75 12.35
16 36.80 37.20 71.52 72.47 11.36 12.63
Qwen-1.5B 2 19.20 24.80 26.60 21.15 24.46 27.56 0.54 0.88 4.22
4 28.60 29.80 25.93 30.09 1.29 4.31
8 33.40 34.20 26.93 28.13 1.93 5.11
16 35.00 36.60 34.27 35.12 4.89 6.96
Llama3-3B 2 31.80 37.00 40.20 73.09 80.29 81.11 14.47 18.01 24.67
4 39.40 41.60 84.15 85.33 21.44 25.51
8 42.60 44.80 85.14 85.78 21.65 26.16
16 44.00 46.40 85.90 88.44 22.94 27.59
Table 2 presents a comparison between our proposed VDS-TTT method and the very recently published TTRL method that uses RL for TTT Zuo et al. (2025). Results show that VDS-TTT outperforms TTRL on the instruction-tuned LLaMA-3.1-8B-Instruct model. This result is particularly striking given that, as acknowledged by TTRL’s authors, reinforcement-learning-based strategies can suffer from sensitivity to query difficulty level, heavy reliance on strong pretraining on similar distributions, and potential training collapse under some conditions. As shown on the challenging AIME2024 dataset with LLaMA-3.1-8B-Instruct, where the model is not specifically tuned for math, TTRL has no effect during test time, while VDS-TTT achieves a 6.7% improvement. Taken together, these findings underscore the practical advantages of VDS-TTT: it achieves competitive and even superior test time performance, is straightforward to implement, and entails significantly less computational overhead compared to the RL-based approaches.
Figure 2 plots the test-time training loss incurred by VDS-TTT for three representative model–benchmark–sampling configurations. In all cases, the loss curves exhibit a smooth, monotonic decrease and plateau at low values, confirming that VDS-TTT consistently adapts model parameters to the test-time distribution and achieves reliable convergence under diverse settings.
(a) Llama-3.2-1B-Instruct on Math-500 for $N=4$
(b) Llama-3.2-3B-Instruct on GSM-8K for $N=2$
(c) DeepSeek-R1-Distill-Qwen-1.5B on AIME for $N=8$
Figure 2: Three instances of TTT loss curves
Figure 3 illustrates the performance trajectory of the iterative VDS-TTT strategy, where the VDS-TTT procedure is applied multiple iterations to the same model. In this figure, the x-axis denotes the number of iterations (i.e., how many times VDS-TTT has been applied), while the y-axis indicates the corresponding accuracy at each iteration. This iterative process reveals a clear trend: model performance consistently improves with successive iterations—up to a certain point. As shown in Table 1 (AIME task), even a single round of VDS-TTT (iteration 1) enables the model to acquire task-specific capabilities, despite having no prior exposure. This motivated us to investigate whether further iterations could yield additional gains. For example, one iteration of VDS-TTT achieves an accuracy of 81.11%, and a second iteration—starting from the fine-tuned checkpoint—leads to further improvements, as the model becomes increasingly capable of generating higher-quality task-relevant outputs. However, due to inherent model capacity limitations, this iterative improvement eventually exhibits diminishing returns. As shown in Figure 3, performance plateaus beyond a certain point, additional iterations result in minor fluctuations rather than continued gains.
The red dashed line in Figure 3 represents the accuracy achieved by the Oracle Verifier. An Oracle Verifier is a theoretical tool that assumes access to ground-truth answers, allowing it to perfectly assess the correctness of model-generated responses. It serves as an upper-bound benchmark to evaluate and compare the effectiveness of practical verification methods used during test time. If an oracle verifier is employed at the first iteration, the model accuracy corresponds to the red dashed line. Remarkably, in several iterations, iterative VDS-TTT surpasses an Oracle verifier demonstrating that self-improvement through multiple iterations can exceed even the ideal “best-of-N” selection at iteration zero. This finding highlights VDS-TTT’s unique ability to not only bootstrap performance from scratch but also to progressively refine model proficiency beyond static verification alone.
A critical insight from Figure 3 is determining the optimal early stopping point for the iterative process. To address this, we define a maximum number of iterations, $T_{\text{max}}$, based on computational budget constraints. We then identify the optimal iteration $t^{*}$, which yields the highest achievable accuracy, using the following criterion: if the difference in accuracy between two consecutive iterations falls below a threshold $\epsilon$ for two successive steps, the process should be halted. Formally, if $A^{(t)}$ denotes the accuracy at iteration $t$, we stop when the accuracy gains between iterations become marginal, indicating convergence. This approach ensures an efficient balance between performance and computational cost by avoiding unnecessary iterations beyond $t^{*}$.
$t^{*}=\min\{t=2,...,T_{max}:|A^{(t)}-A^{(t-1)}|<\epsilon\text{ and }|A^{(t-1)}-A^{(t-2)}|<\epsilon\}$ (2)
(a) Llama-3.2-1B-Instruct on Math-500 for $N=4$
(b) Llama-3.2-3B-Instruct on GSM-8K for $N=2$
Figure 3: Iterative VDS-TTT results
Table 2: Comparison of our proposed framework, VDS-TTT, with the TTRL (Zuo et al. (2025)) approach across three challenging benchmarks. * indicates results from GRPO paper (Hu et al. (2025))
Model Name Methods AIME 2024 AMC Math-500 Average
LLaMA-3.1-8B-Instruct Base model* 3.3 19.3 47.8 23.5
TTRL 3.3 32.5 61.8 32.5
VDS-TTT 10.0 38.5 54.2 38.3
$\Delta$TTRL 0.0 +13.2 +14.0 +9.0
$\Delta$VDS-TTT +6.7 +19.2 +6.4 +10.8
5 Conclusion
We have presented Verifier-Driven sample Selection Test-Time Training (VDS-TTT), a simple yet powerful framework for self-supervised continuous improvement of an LLM model. For each test question, that the LLM responds with high confidence, we choose the single best answer as per a learned verifier to train/adapt the base LLM model on a continuing basis. By using just one response, the test time compute addition is kept low, but the gain in performance is significant, as can be seen from experimental comparisons over other recent methods. By generating $N$ candidate responses per query, scoring them with a verifier, and filtering via a confidence threshold, VDS-TTT constructs a reliable test-time training set on which only low-rank LoRA adapters are updated. This parameter-efficient adaptation mechanism enables rapid convergence and preserves the stability of the base model. Our results confirm that verifier-guided pseudo-labeling addresses the absence of ground-truth labels very well, even under structurally novel reasoning tasks.
Limitations: While VDS-TTT delivers strong adaptation performance on mathematical reasoning benchmarks, its reliance on a verifier trained exclusively on math problems constrains its applicability to other domains. In settings where the underlying verifier has not been exposed to domain-specific reasoning patterns, such as code generation, commonsense QA, or multi-modal tasks, the pseudo-labels it produces may be unreliable, limiting the effectiveness of test-time training. Addressing this limitation will likely require either (1) a general-purpose verifier capable of assessing candidate responses across diverse tasks, or (2) a mixture-of-experts verifier architecture that dynamically activates specialized verification modules based on the input prompt. Such a design would preserve VDS-TTT’s parameter efficiency and test time adaptability while broadening its scope beyond mathematics, thus inviting further exploration into expert routing mechanisms and verifier calibration. Looking forward, we plan to suitably extend VDS-TTT.
Broader Impact: This paper introduces a framework for actively training and adapting a model during test time. By updating on incoming test examples, our approach dynamically aligns the model with new or evolving data distributions in real time and substantially enhances robustness under unpredictable conditions. On the downside, this continual adaptation incurs additional computational overhead, which may pose challenges for large-scale or resource-constrained deployments.
Appendix A Technical Appendices and Supplementary Material
Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline. There is no page limit for the technical appendices.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See Abstract.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Section 5
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: See Section 4.3
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See Section 4.2
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: Publicly available dataset has been used. Code will be released upon publication or provided upon request.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Section 4.2
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: No further statistical experiments were deemed necessary, as the reported results were obtained through multiple runs and consistently fell within the same range, demonstrating their stability and reliability.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Section 4.2
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we confirm
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See Section 5
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: There are no such problems in our task.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: See Section 4.1
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We use existing public benchmark datasets for experiments.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No such experiments or research are involved in our work
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No such experiments or research are involved in our work.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: It hasn’t been used as an important, original, or non-standard component of the core method.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.